OpenStudy (anonymous):

given that x1, x2, x3,......x16 are positive real numbers such that \[\frac{ x1 }{ x2 }=\frac{ x2 }{ x3 }=\frac{ x3 }{ x4 }=....\frac{ x15 }{ x16}\] if x1+x2+x3+x4=20 x5+x6+x7+x8=320, then what is x13+x14+x15+x16 ?

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

this is not so difficult let x1/x2 = ..... = r

OpenStudy (anonymous):

ohk ..,,den .?

hartnn (hartnn):

so, x1 = r x2 = r^2 x3 = r^3 x4 = ....and so on so from x1+x2+x3+x4=20 and x5+x6+x7+x8=320 you should be able to find 'r' try it.

OpenStudy (anonymous):

hm but how r^2 and r^3 came ..?

hartnn (hartnn):

x2/x3 = r x2 = r x3 x1 = r x2 = r (r x3) = r^2 x3

OpenStudy (anonymous):

got got it ! r yes

hartnn (hartnn):

x1+x2+x3+x4=20 r^4 x5 + r^3 x5 + r^2 x5 + r x5 = 20 x5 (r^4+r^3+r^2+r) = 20

OpenStudy (anonymous):

umm ..i did this like \[x1(1+\frac{ 1 }{ r }+\frac{ 1 }{ r^2 }+\frac{ 1 }{ r^3 } ) = 20\]

OpenStudy (anonymous):

something like G.P

hartnn (hartnn):

no prob :) see whether you get r as 1/2 or not.

hartnn (hartnn):

they are surely in GP.

OpenStudy (anonymous):

yes ./.. wow ..r=1/2 ..got that

hartnn (hartnn):

now find x1 from your last equation.

hartnn (hartnn):

then you can find any term, x1 = r x2 = r^2 x3 = r^3 x4 = ....and so on x13 = r^12 x1 x14 = r^13 x1 x15 = r^14 x1 x16 = r^15 x1

OpenStudy (anonymous):

and then again using ..that by converting and putting x1 and r ..in what we need to find .!

OpenStudy (anonymous):

x1 comes out to be 5/3 .??

hartnn (hartnn):

4/3

hartnn (hartnn):

x1 (1+2+4+8) =20 x1 *15 = 20

OpenStudy (anonymous):

hahahah ..lol ...20/15 i wrote ..5/3 ..mad i am

OpenStudy (anonymous):

answer is absolutely ..correct ..thank u

hartnn (hartnn):

welcome ^_^